A 5.0 oz sample of ground beef contains 485 kilocalories and is 21% protein by mass. What is the percent fat, by mass, in the ground beef? Assume all of the kilocalories come from protein and fat.
Answer #1

Sol.

AsTotal Kilocalories of ground beef sample = 485 Kcal

Total Mass of sample = 5 oz = 5 × 28.35 g = 141.75 g  

Also , sample conatins 21 % protein by mass

So , Mass of protein in sample  

= 21 % of Total Mass of sample

= 21 × 141.75 g / 100  

= 29.7675 g

As Calorific value of protein = 4 Kcal / g  

So , Kilocalories come from protein

= 29.7675 g × 4Kcal / g = 119.07 Kcal  

Therefore , Kilocalories come from fat  

= Total Kilocalories of sample - Kilocalories come from protein

= 485 Kcal - 119.07 Kcal  

= 365.93 Kcal  

Now , Calorific value of fat = 9 Kcal / g  

So , Mass of fat in sample  

= 365.93 Kcal / 9 Kcal / g

= 40.6589 g  

Therefore , Percent fat by mass in the sample

= ( Mass of fat in sample / Total Mass of sample ) × 100

= ( 40.6589 g / 141.75 g ) × 100

=   28.68 %